OpenStudy (anonymous):

mulitply then write in simplest form can not be in decimals square root of 12 * square root of 5

OpenStudy (anonymous):

like this? sqrt(12*5)

OpenStudy (anonymous):

sqrt(12*5) = sqrt(12) sqrt(5) = sqrt(4*3) sqrt(5) = sqrt(4) * sqrt(3) * sqrt(5) make sense so far?

OpenStudy (anonymous):

no it's like \[\sqrt{12} * \]

OpenStudy (anonymous):

\[\sqrt{5}\]

OpenStudy (anonymous):

even better, so same idea, break down the sqrt(12) into it's factors the factors of 12 are 4*3 so, sqrt(12) = sqrt(4*3) separate them =sqrt(4) * sqrt(3) simpilfy = 2 * sqrt(3) put the sqrt(5) back in = 2 * sqrt(3) * sqrt(5)

OpenStudy (anonymous):

wouldn't you mutiply them first then do the breaking down?

OpenStudy (anonymous):

we could, but we'd be repeating steps, i'll show you..

OpenStudy (anonymous):

sqrt(12) * sqrt(5) =sqrt(60) =sqrt(5*12) =sqrt(5*3*4) =sqrt(5) sqrt(3) sqrt(4) see?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

really, what every way makes sense to you, do it that way!